# Absorption and scattering of massless scalar waves by Frolov black holes

### v1

@@ Abstract
We comprehensively investigate the absorption and scattering of massless scalar waves by Frolov black holes, which is a class of regularization of the Reissner–Nordström spacetime. By analyzing the null geodesics, we determine the photon sphere radius and the critical impact parameter, deriving the geometric capture cross section and the classical differential scattering cross section. Utilizing the partial-wave method, we numerically compute the absorption and scattering cross sections across a broad frequency range. Our numerical results show excellent agreement with the low-frequency limit and the high-frequency sinc approximation, as well as with the semiclassical glory approximation. We analyze the dependence of the spectra on the charge and the regularization parameter (Hubble length). Under the horizon-radius normalization, we observe that the total absorption cross section increases with the regularization parameter, a behavior we attribute to the variation of the dimensionless mass. Furthermore, by comparing Frolov, Reissner–Nordström, and Hayward black holes, we demonstrate that their absorption and scattering patterns are nearly indistinguishable when their critical impact parameters or glory impact parameters are matched. This indicates that the photon sphere geometry predominantly governs scalar field interactions, while the detailed core structure plays a secondary role. Our work underscores the potential of wave-based probes to test regular black hole geometries and their observable imprints.
@@ I Introduction
For a minimally coupled massless scalar field, this picture is well understood in benchmark geometries. In the zero-frequency limit the total absorption cross section approaches the area of the event horizon [24, 25], while in the high-frequency regime it oscillates around the geometric capture cross section, with an oscillatory pattern that can be related to the properties of the unstable photon orbit and the associated Regge poles of the scattering matrix [17, 46]. These features have been established in detail for Schwarzschild and Reissner–Nordström black holes through a combination of partial-wave analysis and explicit geodesic control [22, 13]. Full-wave computations of scalar scattering in nontrivial charged backgrounds, together with comparisons to geodesic and Regge-pole/glory approximations, have been carried out for charged dilatonic black holes [28].
@@ II.2 Weak deflection and Classical scattering
$\left.\frac{d\sigma_{\rm sc}}{d\Omega}\right|_{\rm cl}=\frac{b}{\sin\theta}\left|\frac{db}{d\theta}\right|.$ (27)
From Eq.(25) and (27), we conclude that the classical differential scattering cross section for small scattering angles is
$\left.\frac{d\sigma_{\rm sc}}{d\Omega}\right|_{\rm cl}=\frac{16M^{2}}{\Theta^{4}}+\frac{15\pi M^{2}}{4\Theta^{3}}-\frac{3\pi Q^{2}}{4\Theta^{3}}+\mathcal{O}\left(\frac{1}{\Theta}\right)^{2}.$ (28)
It is obvious that the Hubble length $\alpha$ has no effect on the dominant term of classical differential scattering cross section.
@@ III.1 Massless scalar field
It is convenient to introduce the partial-wave $S$-matrix, defined by the ratio of the outgoing and ingoing amplitudes
@@ III.2 Absorption cross sections
In the low-frequency limit only the s-wave survives, then the cross section will approach to the area of BH horizon, this universality first established in Ref. [14] and extended to stationary geometries in Ref. [24]. But in high-frequency regime, capture dominates and the cross section rapidly oscillating around the geometric cross section $\sigma_{\rm geo}$, which we have discussed in Sec. II.1. Here we give the expression of absorption cross section in two asymptotic limits
@@ III.3 Scattering cross section
$g(\theta)=\frac{1}{2i\omega}\sum_{l=0}^{\infty}(2l+1)\left[e^{2i\delta_{l}}-1\right]P_{l}(\cos\theta)$ (40)
@@
in this description the phase shifts defined in Eq. (35) encode the same information as the classical deflection function discussed in Sec. II.
@@ IV.2 Scalar absorption
In Fig. 6, we plot the total absorption cross sections (ACS) of massless scalar waves in the Frolov spacetime. The left panel shows the dependence on the Hubble length $\ell$ for fixed charge $Q$, whereas the right panel displays the dependence on $Q$ at fixed $\ell$, illustrating how the ACS changes as the parameters are varied toward the extremal regime.
Figure 6: Total ACS of massless scalar fields in Frolov spacetime. The numerical results (solid lines) are compared with the high-frequency sinc approximation (dotted lines) and the geometric capture limit (dashed lines). Panels display the dependence on the Hubble length $\ell$ at fixed charge $q=0.5$ (left) and on $q$ at fixed $\ell=0.3$ (right)
As we can see in this figure, the numerical results exhibit the expected asymptotic behaviors at both frequency ends. In the low-frequency limit ($\omega r\to 0$), the cross sections universally converge to the horizon area $\sigma_{\rm abs}\to 4\pi r_{h}^{2}$, while in the high-frequency regime, the spectra oscillate around the geometric cross section $\sigma_{\rm abs}\to\sigma_{\rm geo}=\pi b_{c}^{2}$ with a pattern that is well-described by the semiclassical sinc approximation. Apart from these limits, we observe that the overall amplitude of the ACS typically increases as we increase either $q$ or $\ell$. This trend appears opposite to what is reported for RN and Hayward black holes in the standard $M=1$ normalization, where the total ACS usually decreases as the charge or regularizing parameter grows [13, 16]. However, in our $r_{h}=1$ convention the dimensionless mass $m(q,\ell)$ also increases across the allowed parameter space (see Sec. IV and Fig. 5), so part of the apparent growth of the ACS is simply reflecting the change of the overall mass scale rather than a genuine reversal of the underlying geometric effect.
Then we compare the partial ACSs of scalar wave in Frolov spacetime with the same parameters in Fig. 7 as a complementary view of Fig. 6. These plots show explicitly that the parameter dependence of the total ACS is mainly driven by the lowest modes, in particular the $s$-wave ($l=0$), while higher-$l$ modes provide comparatively smaller influence.
Figure 7: Partial absorption cross sections of massless scalar waves in the Frolov spacetime for the lowest angular-momentum modes $l=0,1,2$. The left panel shows the dependence on the Hubble length $\ell$ for fixed charge $Q=0.5$, whereas the right panel displays the dependence on $Q$ with fixed $\ell=0.3$.
@@ IV.4 Similar features for different Frolov BHs
The intriguing work of de Paula et al. [16] investigated under which circumstances the absorption and scattering spectra of Hayward BHs and RN BHs can become almost indistinguishable.
In the absorption problem, they let the critical impact parameters coincide by adjusting appropriate parameters, $b_{c}^{\rm H}(\alpha)=b_{c}^{\rm RN}(Q)$, and showed that, under this condition, the corresponding total ACS are very similar over the whole frequency range. In the scattering case, they instead matched the impact parameter of the backscattered null geodesics, requiring $b_{g}^{\rm H}(\alpha)=b_{g}^{\rm RN}(Q)$, and found that the DCS patterns of two black holes are also nearly identical.
Motivated by this observation, we now ask whether a similar correspondence persists when comparing RN, Hayward and Frolov black holes with the same impact parameters. Since the Frolov spacetime reduces to the RN solution and Hayward black hole in appropriate limits, curves of constant impact parameters $b(\ell,q)$ in the $(\ell^{2},q^{2})$ plane [cf. Fig. 4] naturally interpolate between these two limiting geometries. Along a given iso-$b$ curve there thus exists a continuous family of Frolov BHs sharing the same impact parameter.
To assess how far the absorption and scattering spectra are influenced by these impact parameters, we focus on the lines with $b_{c}\simeq 3\,r_{h}$ and $b_{g}\simeq 3\,r_{h}$, and select three representative configurations on each line: an Reissner–Nordström black hole, a Hayward black hole and an intermediate Frolov case. As shown in Fig. 10(a), we plot the ACSs for three cases with $b_{c}\simeq 3\,r_{h}$ over the whole frequency range. Moreover, Fig. 10(b) illustrates the comparison between the DCSs in the background of three different spacetimes.
(a) The total absorption cross section (ACS)
(b) The differential scattering cross section (DCS)
Figure 10: ACSs and DCSs of Reissner–Nordström, Hayward and Frolov black holes with matched impact parameters. Panel (a) shows configurations with critical impact parameter $b_{c}\simeq 3\,r_{h}$, whereas panel (b) refers to configurations with glory impact parameter $b_{g}\simeq 3\,r_{h}$.
We exhibit, in both panels of Fig. 10, that the ACSs and DCSs for three chosen black hole are basically coincide in the whole frequency/angle range. While the agreement in the asymptotic limits is guaranteed by our parameter matching strategy (fixing $r_{h}$ for the low-frequency limit and $b_{c}/b_{g}$ for the high-frequency limit), the nearly perfect overlap in the intermediate regime is non-trivial. This implies that the iso-spectrality is not merely an artifact of the limits, but stems from the fact that the associated effective potentials possess almost identical barrier shapes around the photon sphere. Consequently, the Frolov, Reissner–Nordström, and Hayward black holes become practically indistinguishable via massless scalar scattering in these configurations.
@@ V Conclusion
In the numerical analysis we adopted the normalization $r_{h}=1$ and mapped the allowed region in the $(\ell^{2},q^{2})$ plane where an event horizon exists. In this parametrization the normalized mass $m$ grows monotonically with both $q$ and $\ell$, so changing these parameters at fixed $r_{h}$ does not correspond to the same effect on spacetime as in the more common $M=1$ convention. This explains why, in our results, the total absorption cross section and the height of the effective potential barrier increase as either parameter is raised, even though related studies of Reissner–Nordström and Hayward black holes at fixed mass report a decreasing trend. The partial–wave decomposition further shows that the lowest mode dominate the low–frequency absorption and drive most of the parameter dependence, while variations of parameters in the scattering problem mainly shift and distort the interference fringe pattern.

### v2

@@ Abstract
We study the absorption and scattering of massless scalar waves by Frolov black holes, a regular deformation of the Reissner–Nordström geometry. A null-geodesic analysis provides the photon-sphere radius and the critical impact parameters governing capture and glory scattering. Using a partial-wave approach, we compute total/partial absorption cross sections and differential scattering cross sections over a broad frequency range. When the absorption spectrum is rescaled by photon-sphere scales, the high-frequency oscillations exhibit a pronounced data collapse in the variables $\hat{\sigma}=\sigma_{\rm abs}/\sigma_{\rm geo}$ and $x=\omega/\Omega_{c}$, highlighting photon-sphere control of the absorption fine structure. We clarify the parameter dependence under the horizon-radius normalization and relate the apparent trends to the variation of the dimensionless mass across parameter space. Finally, comparing Frolov, Reissner–Nordström, and Hayward black holes with matched critical or glory impact parameters, we find that their absorption and scattering patterns can become remarkably close, indicating that in the intermediate-to-high frequency regime the dominant signatures are largely governed by the unstable photon orbit, while core effects enter as subleading corrections.
@@ I Introduction
For a minimally coupled massless scalar field, this picture is well understood in benchmark geometries. In the zero-frequency limit the total absorption cross section approaches the area of the event horizon [24, 25], while in the high-frequency regime it oscillates around the geometric capture cross section, with an oscillatory pattern that can be related to the properties of the unstable photon orbit and the associated Regge poles of the scattering matrix [17, 46]. These features have been established in detail for Schwarzschild and Reissner–Nordström black holes within the partial-wave framework pioneered for black-hole scattering by Sánchez [47] and developed further in subsequent work (see, e.g., [22, 13]). Full-wave computations of scalar scattering in nontrivial charged backgrounds, together with comparisons to geodesic and Regge-pole/glory approximations, have been carried out for charged dilatonic black holes [28].
@@ II.2 Weak deflection and Classical scattering
$\left.\frac{d\sigma_{\rm sc}}{d\Omega}\right|_{\rm cl}=\frac{b(\Theta)}{\sin\theta}\left|\frac{db(\Theta)}{d\Theta}\right|,$ (27)
where $\theta$ is the scattering angle, and $\Theta$ is the deflection angle. From Eq.(25) and (27), we conclude that the classical differential scattering cross section for small scattering angles is
$\left.\frac{d\sigma_{\rm sc}}{d\Omega}\right|_{\rm cl}=\frac{16M^{2}}{\Theta^{4}}+\frac{3\pi(5M^{2}-Q^{2})}{4\Theta^{3}}+\mathcal{O}(\Theta^{-2}).$ (28)
Eq. (28) immediately connects with known results for other black-hole spacetimes. For $Q=0$ it reduces to the Schwarzschild small-angle cross section, whose universal leading term $16M^{2}/\Theta^{4}$ corresponds to gravitational Rutherford scattering, together with the standard correction $\propto\Theta^{-3}$ [47, 44]. For $\alpha=0$ the expression coincides with the Reissner–Nordström case, where the electric charge first enters at order $\Theta^{-3}$ and decreases the cross section relative to Schwarzschild [13]. In the Frolov spacetime the parameter $\alpha$ does not contribute up to order $\Theta^{-3}$, this follows from the large-$r$ expansion of f(r)f(r)(see Eq. (26)), where the $\alpha$-dependent terms start at $\mathcal{O}(r^{-4})$, so that they affect the weak-deflection series only at higher orders, as also observed in the Hayward case [16].
@@ III.1 Massless scalar field
Following the standard partial-wave treatment of wave scattering and absorption by spherically symmetric black holes [47, 46, 22, 13], we impose purely ingoing behavior at the horizon and an incident-plus-outgoing decomposition at infinity, which leads to the boundary conditions in Eq. (34) and the associated flux relation. It is convenient to introduce the partial-wave $S$-matrix, defined by the ratio of the outgoing and ingoing amplitudes
@@ III.2 Absorption cross sections
In the low-frequency limit only the s-wave survives, then the cross section approaches the area of BH horizon, this universality first established in Ref. [14] and extended to stationary geometries in Ref. [24]. But in the high-frequency regime, capture dominates and the cross section rapidly oscillates around the geometric cross section $\sigma_{\rm geo}$, as we have discussed in Sec. II.1. Here we give the expression of absorption cross section in two asymptotic limits
@@ III.3 Scattering cross section
$g(\theta)=\frac{1}{2i\omega}\sum_{l=0}^{\infty}(2l+1)\left[S_{l}(\omega)-1\right]P_{l}(\cos\theta)$ (40)
@@
In this description the complex quantity $S_{l}(\omega)$ encodes the same physical information as the classical deflection function discussed in Sec. II.
@@ IV.2 Scalar absorption
In Fig. 6, we present the total absorption cross section $\sigma_{\rm abs}/(\pi r_{h}^{2})$ of massless scalar waves for representative values of $(q,\ell)$. For reference we also display the geometric-optics limit $\sigma_{\rm geo}=\pi b_{c}^{2}$ and the high-frequency approximation $\sigma_{\rm hf}$ in Eq. (39).
Figure 6: Total absorption cross section $\sigma_{\rm abs}$ of massless scalar waves by Frolov black holes, normalized by $\pi r_{h}^{2}$, as a function of $\omega r_{h}$. In the left panel we fix $q=0.5$ and vary $\ell$, while in the right panel we fix $\ell=0.3$ and vary $q$. Solid curves show the numerical partial-wave result. The horizontal dashed lines indicate the geometric-optics cross section $\sigma_{\rm geo}=\pi b_{c}^{2}$ determined by the critical null geodesic. The dotted curves correspond to the high-frequency approximation $\sigma_{\rm hf}$ in Eq. (39), which reproduces the high-frequency oscillations of $\sigma_{\rm abs}$ around $\sigma_{\rm geo}$ with the universal Regge-pole/sinc structure controlled by photon-sphere parameters (e.g. $\Omega_{c}$ and the Lyapunov exponent) [17]. In the low-frequency limit, $\sigma_{\rm abs}\to A_{h}=4\pi r_{h}^{2}$ as in Ref. [14, 24, 46].
We observe that the numerical results exhibit the standard low- and high-frequency behaviors as expected from the general theory of black-hole absorption. We also note that the parameter dependence of $\sigma_{\rm abs}/(\pi r_{h}^{2})$ under the $r_{h}=1$ normalization typically increases when either $q$ or $\ell$ is increased within the allowed domain. This trend should be interpreted with some care when compared with the common $M=1$ normalization used for RN/Hayward black holes, for which the total ACS often decreases as the charge or Hubble length grows [13, 16]. In our convention the dimensionless mass $m(q,\ell)$ varies across parameter space (see Fig. 5), so part of the apparent increase of $\sigma_{\rm abs}/(\pi r_{h}^{2})$ reflects the change of the overall mass scale, in addition to the genuine variation of the photon-sphere data entering $b_{c}/r_{h}$.
To disentangle the nontrivial wave effects from the overall geometric scaling induced by varying $(q,\ell)$ at fixed $r_{h}=1$, it is useful to re-express the absorption spectrum in terms of photon-sphere scales. We therefore consider the rescaled quantities $\hat{\sigma}\equiv\sigma_{\rm abs}/\sigma_{\rm geo}$ and $x\equiv\omega/\Omega_{c}$. For a static spherically symmetric geometry one has $\Omega_{c}=\sqrt{f(r_{c})}/r_{c}=1/b_{c}$, so $x=\omega b_{c}$. Since $\hat{\sigma}$ and $x$ are dimensionless and invariant under an overall length rescaling, this representation does not depend on whether one adopts $r_{h}$ or $M$ as the fundamental unit. As shown in Fig. 7, the high-frequency oscillations for different $(q,\ell)$ largely collapse in these variables, supporting the interpretation that the dominant high-frequency structure is controlled by photon-sphere data ($b_{c}$, $\Omega_{c}$ and the Lyapunov exponent entering the Regge-pole/sinc formula), while residual differences at intermediate frequencies reflect subleading (beyond-eikonal) effects.
Figure 7: Scaled total absorption cross section shown as $\sigma_{\rm abs}/\sigma_{\rm geo}$ versus $\omega/\Omega_{c}$ (equivalently $\omega b_{c}$ for static spherical spacetimes). The left panel fixes $q=0.5$ and varies $\ell$, while the right panel fixes $\ell=0.3$ and varies $q$, matching the parameter slices of Fig. 6. The collapse of the high-frequency oscillations in these rescaled variables highlights the photon-sphere dominance of the absorption fine structure (cf. the universal Regge-pole/sinc description) [17].
As a complementary view, Fig. 8 shows the partial absorption cross sections $\sigma_{l}$ for $l=0,1,2$ for the same parameter choices. The total ACS trend is largely controlled by the lowest modes: at low frequencies the $s$-wave dominates, while higher-$l$ channels are progressively activated only when the wave can efficiently penetrate the corresponding effective-potential barrier.
Figure 8: Partial absorption cross sections $\sigma_{l}$ for the lowest modes $l=0,1,2$, normalized by $\pi r_{h}^{2}$, as functions of $\omega r_{h}$. In the low-frequency limit, the absorption is dominated by the $s$-wave and $\sigma_{\rm abs}\to A_{h}$, whereas higher modes are suppressed by the centrifugal barrier ($\sigma_{l>0}\to 0$ as $\omega\to 0$) [46, 24]. As $\omega$ increases, each partial wave is activated when the wave can penetrate the corresponding effective-potential barrier, leading to the characteristic rise and peak structure familiar from Schwarzschild/RN and other (regular) black holes [22, 13].
@@ IV.4 Similar features for different Frolov BHs
The intriguing work of de Paula et al. [16] investigated under which circumstances the absorption and scattering spectra of Hayward and Reissner–Nordström (RN) black holes can become nearly indistinguishable. In the absorption problem, they matched the critical impact parameters by tuning the model parameters, $b_{c}^{\rm H}(\alpha)=b_{c}^{\rm RN}(Q)$, and found that the corresponding total absorption cross sections (ACSs) are very similar over a broad frequency range. In the scattering problem, they instead matched the impact parameter of the backscattered null geodesics, $b_{g}^{\rm H}(\alpha)=b_{g}^{\rm RN}(Q)$, and observed that the differential scattering cross sections (DCSs) exhibit nearly identical interference patterns.
Motivated by this observation, we ask whether a similar correspondence persists when comparing RN, Hayward, and Frolov black holes. Since the Frolov spacetime reduces to RN and Hayward geometries in appropriate limits, curves of constant impact parameters $b(\ell,q)$ in the $(\ell^{2},q^{2})$ plane (cf. Fig. 4) naturally interpolate between these limiting cases. Along a given iso-$b$ curve there therefore exists a continuous family of Frolov black holes sharing the same characteristic impact parameter.
To assess the extent to which the wave observables are controlled by these impact parameters,we focus on the representative lines $b_{c}\simeq 3\,r_{h}$ and $b_{g}\simeq 3\,r_{h}$ and select three configurations on each line: an RN black hole, a Hayward black hole, and an intermediate Frolov case. Figure 11 compares the corresponding ACSs (left panel, matched $b_{c}$) and DCSs (right panel, matched $b_{g}$).
Figure 11: ACSs and DCSs of Reissner–Nordström, Hayward and Frolov black holes with matched impact parameters. The left panel compares ACSs for configurations with $b_{c}\simeq 3\,r_{h}$, while the right panel compares DCSs for configurations with $b_{g}\simeq 3\,r_{h}$.
We observe that, under this matching strategy, the ACSs and DCSs of the three geometries nearly coincide over the displayed frequency/angle ranges. While the agreement in the asymptotic limits is expected (the low-frequency behavior is governed by $r_{h}$, whereas the high-frequency structure is controlled by the photon-sphere data encoded in $b_{c}$ or $b_{g}$), the close overlap in the intermediate regime is nontrivial. This suggests that the relevant effective potentials possess very similar barrier shapes in the vicinity of the photon sphere, so that the dominant features of massless-scalar absorption and scattering are largely controlled by the unstable null orbit, with the detailed regular-core structure contributing only subleading corrections in these configurations. Accordingly, within the present probe (massless scalar) and the parameter-matched setup, distinguishing the three spacetimes using these wave observables alone can be challenging.
@@ V Conclusion
In the numerical analysis we adopted the normalization $r_{h}=1$ and mapped the allowed region in the $(\ell^{2},q^{2})$ plane where an event horizon exists. In this parametrization the normalized mass $m$ grows monotonically with both $q$ and $\ell$, so changing these parameters at fixed $r_{h}$ does not correspond to the same physical variation as in the more common $M=1$ convention. This explains why, in our results, the total absorption cross section and the height of the effective potential barrier increase as either parameter is raised, even though related studies of Reissner–Nordström and Hayward black holes at fixed mass report a decreasing trend. The partial-wave decomposition further shows that the lowest mode dominates the low-frequency absorption and drives most of the parameter dependence, while varying $(q,\ell)$ in the scattering problem mainly shifts and distorts the interference fringe pattern.
We also presented a scale-independent characterization of the absorption spectrum by rescaling the results with photon-sphere scales. Using $\hat{\sigma}=\sigma_{\rm abs}/\sigma_{\rm geo}$ and $x=\omega/\Omega_{c}$, we found that the high-frequency oscillations for different $(q,\ell)$ largely collapse onto a nearly universal curve, see Fig. 7. Because both $\hat{\sigma}$ and $x$ are dimensionless and invariant under an overall length rescaling, this collapse is insensitive to whether one adopts $r_{h}$ or $M$ as the fundamental unit. It highlights that the dominant high-frequency structure is primarily encoded in photon-sphere data, while residual differences at intermediate frequencies reflect subleading effects beyond the eikonal description.